OpenStudy (anonymous):

Can someone show me how to turn standard form equations into slope intercept form equations. Example: 2x=5y=12 I GIVE MEDALS!!!

OpenStudy (anonymous):

too many equal signs

OpenStudy (anonymous):

i think the question is wrong it should have been 2x+5y=12 or maybe 2x-5y=12

OpenStudy (anonymous):

2x+5y=12*

OpenStudy (anonymous):

now if its 2x+5y=12 divide by 5 y=-(2/5)x+12/5 this is the required form slope = (-2/5) and intercept = 12/5

OpenStudy (anonymous):

ok thank you